# Nightjar: Dynamic Adaptive Speculative Decoding for Large Language Models Serving

### v2

@@ 5   Receive current batch size $B=B_{t}$
7    // Bin Initialization
8      Set bin type: Exploration with prob. $1/\sqrt{b_{B}}$, else Exploitation
9 if bin is Exploration then
10    // Uniform Exploration
11      Sample $\gamma_{t}\sim\text{Uniform}(\{0,\dots,\Gamma_{\max}\})$
@@
13    // Cost-Aware Exploitation
14      Select $\gamma_{t}$ via Eq. (3)
15  Play $\gamma_{t}$, observe reward $r_{t}(\gamma_{t})$ and update $\tilde{g}_{B,\gamma}$
@@ 24 $t\leftarrow t+1$
Overhead. The arm selection execution time is approximately 1e-5 seconds, while the average generation time per token in the LLM is 0.034 seconds — a difference of nearly 3400 times. This minimal overhead demonstrates the exceptional efficiency of our method, as the cost of dynamic arm selection is negligible compared to the intrinsic latency of token generation.
4.3 Theoretical Analysis
In this section, we provide a rigorous theoretical analysis for our proposed Nightjar algorithm. We begin by presenting our main theoretical result, Theorem 1, which establishes a sublinear regret bound for the algorithm under the dynamic serving environment. The proof relies on decomposing the regret into exploration, estimation, and switching cost components.
For the analysis, we assume the goodput reward $r_{t}$ is bounded in $0,g_{\max}]$ and the switching cost is bounded by $C_{\text{switch}}$.
4.3.1 Regret of Nightjar
The core element of our proposed framework is the adaptive bin-greedy strategy. The following theorem presents the regret upper bound of Nightjar.
Theorem 1.
For any time horizon $T$ and batch size context $B$, assuming the switching cost is bounded, there exist positive constants such that with probability at least $1-\delta$, the cumulative regret of the proposed Nightjar algorithm satisfies:
$R(T)\leq\mathcal{O}\left(g_{\max}\sqrt{T\log T}+C_{\text{switch}}\sqrt{T}\right)=\tilde{\mathcal{O}}(\sqrt{T}).$ (4)
Consequently, the average regret satisfies $\lim_{T\to\infty}R(T)/T=0$.
Remark 1.
Theorem 1 provides a high-probability theoretical guarantee. Unlike expected regret bounds, this ensures that the algorithm performs efficiently in widely varying single runs, which is critical for system stability in production environments.
4.3.2 Proof of Theorem 1
To prove Theorem 1, we introduce three supporting lemmas handling estimation confidence, exploration frequency, and switching frequency.
Lemma 1 (Confidence Bound).
For any arm $\gamma$ and time $t$, let $N_{t}(\gamma)$ be the number of times arm $\gamma$ has been pulled. With probability at least $1-\delta/2$, the deviation between the empirical mean goodput $\tilde{g}_{t,\gamma}$ and the true mean $\mu_{\gamma}$ is bounded by the confidence radius $\beta_{t}(\gamma)$:
$|\tilde{g}_{t,\gamma}-\mu_{\gamma}|\leq\beta_{t}(\gamma)\coloneqq\sqrt{\frac{g_{\max}^{2}\log(2T/\delta)}{2N_{t}(\gamma)}}.$ (5)
Lemma 2 (Bounding Exploration).
Under the adaptive binning mechanism, the exploration probability in the $b$-th bin is set exactly to $p_{b}=1/\sqrt{b}$. The cumulative number of exploration steps up to time $T$ is bounded by:
$N_{\text{explr}}(T)=\sum_{t=1}^{T}\mathbb{I}(\text{exploration}_{t})\leq\mathcal{O}(\sqrt{T}).$ (6)
Lemma 3 (Bounding Switching Cost).
The total number of speculative length switches, denoted as $S(T)$, is bounded by the sum of exploration switches and exploitation updates. Given the block-bin structure where block size grows geometrically, $S(T)$ satisfies:
$S(T)\leq\mathcal{O}(\sqrt{T}).$ (7)
Indeed, switches primarily occur during exploration steps (bounded by Lemma 2) or at bin boundaries; since bin sizes grow with $\sqrt{H_{B}}$, the frequency of boundary updates decays, which yields $S(T)\leq\mathcal{O}(\sqrt{T})$.
Based on Lemmas 1, 2, and 3, we derive Theorem 1.
Proof of Theorem 1.
The cumulative regret $R(T)$ can be decomposed into three parts: regret from exploration rounds ($R_{\text{explr}}$), regret from suboptimal choices during exploitation due to estimation error ($R_{\text{est}}$), and the cumulative switching cost ($C_{\text{switch}}\cdot S(T)$).
$R(T)$ $=R_{\text{explr}}+R_{\text{est}}+\text{SwitchingCost}(T)$
$\leq\underbrace{N_{\text{explr}}(T)\cdot g_{\max}}_{(a)}+\underbrace{\sum_{t\in\mathcal{T}_{\text{exploit}}}(\mu^{*}-\mu_{\gamma_{t}})}_{(b)}+\underbrace{S(T)\cdot C_{\text{switch}}}_{(c)}$
Term (a): By Lemma 2, the exploration regret is bounded by $g_{\max}\cdot\mathcal{O}(\sqrt{T})$.
Term (c): By Lemma 3, the switching cost is bounded by $C_{\text{switch}}\cdot\mathcal{O}(\sqrt{T})$.
Term (b): For exploitation steps, if a suboptimal arm $\gamma_{t}$ is chosen, it implies the empirical estimate $\tilde{g}_{t,\gamma_{t}}$ exceeded the optimal $\tilde{g}_{t,\gamma^{*}}$. Using the confidence bound from Lemma 1, the instantaneous regret is bounded by $2\beta_{t}(\gamma_{t})$. Summing over $T$, standard bandit analysis (deriving from the summation of inverse square roots) yields:
$\sum_{t=1}^{T}2\sqrt{\frac{g_{\max}^{2}\log(2T/\delta)}{2N_{t}(\gamma_{t})}}\leq\mathcal{O}\left(g_{\max}\sqrt{T\log T}\right).$ (8)
Combining these terms:
$R(T)$ $\leq\mathcal{O}(g_{\max}\sqrt{T})+\mathcal{O}(g_{\max}\sqrt{T\log T})+\mathcal{O}(C_{\text{switch}}\sqrt{T})$
$=\mathcal{O}\left(\sqrt{T\log T}\right).$ (9)
This completes the proof. ∎
@@ 5 Elastic Memory Management for Model-Based Draft Models
Under high system contention, characterized by a backlog in the request queue, the scarcity of KV cache memory often necessitates frequent swap operations or the rejection of active requests. In such saturation regimes, speculative decoding becomes counter-productive due to scheduling overheads. To address this, we propose a dynamic offloading mechanism for model-based draft models. As illustrated in Figure 6, we offload the draft model weights—analogous to LoRA adapter management—to the host memory, thereby reclaiming GPU memory to expand the capacity of the original KV cache. Let $M_{\text{orig}}$ denote the baseline KV cache capacity; the expanded capacity is defined as $M_{\text{scale}}=M_{\text{orig}}+S_{\text{draft}}$, where $S_{\text{draft}}$ represents the memory footprint of the draft model.
@@ 5.1 Trigger Conditions for Dynamic Adjustment
Conversely, a KV cache contraction is initiated to restore the draft model when: (1) The waiting queue is empty ($|\mathcal{Q}_{\text{wait}}|=0$); (2) Sufficient memory exists to accommodate both the draft model and a safety buffer ($N_{\text{free}}>S_{\text{draft}}+\tau_{\text{low}}$). This hysteresis design ensures that speculative decoding remains disabled during high-load phases (where $N_{\text{free}}<\tau_{\text{low}}$), thereby preventing resource thrashing.
@@ 5.2 Mechanism of KV Cache Expansion
Upon triggering expansion, the system dynamically reallocates GPU memory. The Model Executor allocates a contiguous memory segment of size $S_{\text{draft}}$ to the KV cache pool. Subsequently, the Block Manager updates its internal state: (1) The set of allocatable physical block indices is expanded to include the new range $K_{\text{curr}},K_{\text{new}})$; (2) Reference counters for these newly instantiated blocks are initialized to zero; (3) The new block indices are appended to the free block queue. Crucially, this process is implemented via in-place expansion, eliminating the need for massive memory copying and minimizing reconfiguration overhead.
@@ 5.3 KV Cache Contraction and Logical Remapping
Figure 7: Overview of the KV Cache Block Contraction and Logical Table Remapping process. The system ensures data consistency during physical memory compaction through topological sorting and vectorized data migration.
@@
Step 1: Identification of Eviction Candidates. The system scans the page tables of all active sequences to identify the set of physical blocks $\mathcal{B}_{\text{evict}}$ located in the memory region to be reclaimed (i.e., blocks with $\text{ID}\geq K_{\text{new}}$). These blocks must be migrated to the preserved memory region $0,K_{\text{new}})$.
Step 2: Construction of Migration Plan. We verify available slots in the free list with IDs strictly less than $K_{\text{new}}$. A mapping $\mathcal{M}:b_{\text{old}}\rightarrow b_{\text{new}}$ is constructed. To preserve logical dependencies (specifically, the pointers to predecessor blocks), we apply a topological sort to the migration set. This ensures that blocks are migrated in an order that maintains the integrity of the dependency chain.
Step 3: Metadata Update and Remapping. For each block requiring migration, the system performs atomic metadata updates: (a) A new physical block object is instantiated, inheriting the token sequence and size of the original; (b) Predecessor pointers are updated—if a predecessor block is also part of the migration set, the pointer is redirected to its new location; otherwise, the original reference is maintained; (c) The logical Block Table is updated to map the sequence to the new physical block ID; (d) The reference counter for the new block is incremented, and it is marked as allocated. Finally, the old physical block is released, ensuring the logical sequence remains consistent with the physical memory state.
Step 4: Vectorized Data Migration. Following the metadata update, the physical data transfer is executed. To handle potential circular dependencies in the mapping (e.g., $A\to B$ and $B\to A$), the topological sort resolves execution order, isolating any simplified cyclic swaps. The migration is powered by a custom high-performance kernel (implemented in Triton) utilizing a tile-based parallelization strategy. Each thread block manages the transfer of a complete KV cache block via vectorized load/store instructions. This approach provides robust performance even on hardware lacking advanced virtual memory management (VMM) features like cuMemMap. By decoupling memory reclamation from model execution, we eliminate deadlock risks associated with ”hold-and-wait” patterns.
Step 5: Allocator State Finalization. Post-migration, the block allocator trims its index set, removing IDs $\geq K_{\text{new}}$ and clearing associated reference counts, effectively finalizing the contraction.
@@ 5.4 Non-blocking Asynchronous Migration
To minimize latency, we leverage the fact that speculative decoding is inactive during the unloading phase. Consequently, draft model weights are evicted from the GPU immediately, with a backup retained in host memory. Conversely, when VRAM availability indicates a low request rate, the draft model is reloaded to resume accelerated generation. This reloading process utilizes asynchronous CUDA streams with non-blocking Direct Memory Access (DMA). This allows data transfer to overlap with ongoing computation, ensuring that the background management of model weights incurs negligible impact on the serving latency of the primary model.
@@ 6.2 Dynamic request load
For 7B model under dynamic request rate (Figure 10), as shown in Figure 11, Nightjar surges ahead during low request load periods, achieving the highest peak throughput ( 3800 tokens/s) and maintaining a consistent lead over rivals during high load phases. DSD’s predictions are less accurate during low request load, leading to throughput degradation, but perform better under high request load, while TETRIS exhibits performance drops under high request load due to its high overhead in large batch size and static BanditSpec is not suitable for dynamic batch size.
@@
As shown in Table IV, Nightjar consistently achieves the highest throughput, demonstrating superior scalability over both static and dynamic baselines. While SD provides moderate gains, Nightjar significantly amplifies these benefits. Notably, on the 13B model, Nightjar achieves a remarkable 64.6% improvement over vanilla decoding (w/o SD) on the Alpaca dataset. More importantly, Nightjar shows exceptional robustness in complex scenarios. Compared to the advanced inference system TETRIS, Nightjar maintains a consistent lead, achieving a 12.1% throughput gain on the 7B Alpaca dataset. Furthermore, on the 13B SpecBench dataset, Nightjar outperforms the competing DSD method by a substantial 53.4% and BanditSpec by 39.1%. Even against the optimized fixed-length SD baseline, Nightjar maintains a clear advantage, proving that its cost-aware policy effectively identifies optimal lengths.
Regarding end-to-end latency, Table V further validates Nightjar’s efficiency. For the 7B model, Nightjar reduces latency by 20.2% compared to SD on the Alpaca dataset. Crucially, Nightjar demonstrates superior lightweight scheduling compared to TETRIS. On the 7B Alpaca dataset, Nightjar reduces latency by 24.4% compared to TETRIS, indicating that Nightjar effectively mitigates the complex sampling overheads that hinder TETRIS. On the 13B model, the advantage remains significant against other dynamic baselines: Nightjar reduces latency by 34.3% compared to DSD on SpecBench. This confirms that Nightjar’s bandit-based mechanism avoids the ”decision deadlock” and high penalty costs that hamper DSD in high-load, compute-bound environments.

### v3

@@ 5   Receive current batch size $B=B_{t}$
7    // Bin Start: Select Strategy & Arm
8      Let $p=1/b_{B}$
9      Sample $u\sim\text{Uniform}(0,1)$
10    if $u<p$ then
11       // Exploration: Random Arm
12       $\gamma_{\text{curr}}\sim\text{Uniform}(\{0,\dots,\Gamma_{\max}\})$
@@
14       // Exploitation: Best Arm (Lock for Bin)
[…]
16 // Execute the locked arm for this step
17   Play $\gamma_{t}\leftarrow\gamma_{\text{curr}}$
18   Observe goodput $g_{t}$ and update $\tilde{g}_{B,\gamma_{t}}$
@@ 27 $t\leftarrow t+1$
Overhead. The arm selection execution time is approximately 1e-5 seconds, while the average generation time per token in the LLM is 0.034 seconds, a difference of nearly 3400 times. This minimal overhead demonstrates the exceptional efficiency of our method, as the cost of dynamic arm selection is negligible compared to the intrinsic latency of token generation.
5 Theoretical Analysis of Regret Bound
In this section, we provide a theoretical analysis of the cumulative regret for the proposed Nightjar algorithm. We demonstrate that the hierarchical bin-based structure effectively controls the switching cost, ensuring sublinear regret.
5.1 Preliminaries and Loss Formulation
Unlike standard Multi-Armed Bandit (MAB) settings that maximize reward, Nightjar minimizes the latency per token. Let $\ell_{B,\gamma}=1/g_{B,\gamma}$ denote the expected latency per token for a given batch size $B$ and speculative length $\gamma$. We define the switching cost (KV cache reconstruction overhead) as $C_{\text{switch}}$.
We formulate the Loss at time step $t$ as:
$L_{t}(\gamma_{t})=\ell_{B_{t},\gamma_{t}}+\mathbb{I}(\gamma_{t-1}=0\land\gamma_{t}>0)\cdot\frac{C_{\text{switch}}}{\gamma_{t}}$ (5)
where the second term represents the switching latency amortized over the generated tokens.
The Cumulative Regret $R(T)$ over a time horizon $T$ is defined as the difference between the total loss incurred by the algorithm and the loss of the optimal fixed policy $\pi^{*}$ (which knows the true distributions):
$R(T)=\sum_{t=1}^{T}\left(L_{t}(\gamma_{t})-L_{t}(\gamma^{*})\right)$ (6)
where $\gamma^{*}=\arg\min_{\gamma}\mathbb{E}[\ell_{B,\gamma}]$.
5.2 Upper Bound on Switching Cost
A critical challenge in dynamic speculative decoding is that frequent changes in $\gamma$ incur high $C_{\text{switch}}$ costs. Nightjar mitigates this via the Bin Locking Mechanism, where the selected arm $\gamma$ is fixed for the duration of a bin.
Lemma 1 (Bound on Number of Switches).
The total number of strategy switches $S_{T}$ up to time $T$ is bounded by the number of bins initiated by the algorithm. Consequently, the cumulative switching cost is bounded by $\mathcal{O}(\sqrt{T})$.
Proof.
In Nightjar, the speculative length $\gamma$ is re-evaluated only at the start of a bin (when $\tau_{B}=1$). Within a bin, the arm remains constant. The algorithm organizes time into blocks indexed by $j$. For a block $j$, the duration is $H_{j}=2^{j-1}$, and the bin size is set to $\sqrt{H_{j}}$. The number of bins in block $j$ is thus $\sqrt{H_{j}}$.
Summing over all blocks up to time $T$, the total number of switches $S_{T}$ is bounded by:
$S_{T}\leq\sum_{j=1}^{\log_{2}T}\sqrt{2^{j-1}}\approx\sum_{j=1}^{\log_{2}T}2^{(j-1)/2}=\mathcal{O}(\sqrt{T})$ (7)
Thus, the total cumulative switching cost is:
$\text{Cost}_{\text{switch}}(T)\leq S_{T}\cdot C_{\max}=\mathcal{O}(C_{\max}\sqrt{T})$ (8)
This ensures that the switching cost component of the regret remains sublinear, preventing the linear regret scenario caused by thrashing. ∎
5.3 Regret Decomposition and Bound
We decompose the remaining regret (excluding switching costs) into exploration and exploitation components. Our analysis adapts the framework of ADA-BINGREEDY.
Theorem 1 (Regret Bound).
Assuming a stationary environment where the latency distributions for each batch size $B$ do not change over time, and by setting the exploration probability $p_{b_{B}}\propto 1/b_{B}$, the Nightjar algorithm achieves a cumulative regret of:
$R(T)=\tilde{\mathcal{O}}(\sqrt{T})$ (9)
The regret is composed of switching costs and sampling regret (exploration and estimation error).
  1. 1.
    Switching Cost: From Lemma 1, the number of switches is bounded by the number of bins. Since the bin size scales with $\sqrt{H_{B}}$, the total switching cost is bounded by $\mathcal{O}(\sqrt{T})$.
  2. 2.
    Exploration Regret: With the updated exploration probability $p=1/b_{B}$, the expected number of exploration bins in a block of duration $H$ (containing $K=\sqrt{H}$ bins) is $\sum_{b=1}^{K}\frac{1}{b}\approx\ln K$. Since each bin has duration $\sqrt{H}$, the regret incurred by random exploration in this block is $\mathcal{O}(\sqrt{H}\ln\sqrt{H})$. Summing over all blocks up to $T$ yields $\sum_{j}\sqrt{2^{j}}j\approx\tilde{\mathcal{O}}(\sqrt{T})$.
  3. 3.
    Exploitation Regret: In exploitation bins, we select the arm with the best empirical regularized latency. Using standard concentration inequalities (e.g., Hoeffding-Azuma), the probability of selecting a suboptimal arm decays exponentially with the number of samples. The analysis follows standard ”Batched Bandit” techniques [9], showing that the exploitation regret is dominated by the exploration and switching terms.
Combining these terms, the total regret is $R(T)=\tilde{\mathcal{O}}(\sqrt{T})$.
We provide a complete proof in Appendix A.
@@ 6 Elastic Memory Management for Model-Based Draft Models
Under high system contention, characterized by a backlog in the request queue, the scarcity of KV cache memory often necessitates frequent swap operations or the rejection of active requests. In such saturation regimes, speculative decoding becomes counter-productive due to scheduling overheads. To address this, we propose a dynamic offloading mechanism for model-based draft models. As illustrated in Figure 6, we offload the draft model weights to the host memory, analogous to LoRA adapter management, thereby reclaiming GPU memory to expand the capacity of the original KV cache. We measure both KV cache and draft model in memory size (bytes) for unit consistency. Let $M_{\text{orig}}$ denote the baseline KV cache memory size (bytes); the expanded memory size is $M_{\text{scale}}=M_{\text{orig}}+S_{\text{draft}}$, where $S_{\text{draft}}$ is the draft model weight memory size (bytes). Equivalently, in block count: $N_{\text{scale}}=N_{\text{orig}}+N_{\text{draft}}$ with $N_{\text{draft}}=\lceil S_{\text{draft}}/B_{\text{block}}\rceil$ and $B_{\text{block}}$ the bytes per KV cache block.
@@ 6.1 Trigger Conditions for Dynamic Adjustment
Conversely, a KV cache contraction is initiated to restore the draft model when: (1) The waiting queue is empty ($|\mathcal{Q}_{\text{wait}}|=0$); (2) Sufficient free blocks exist to accommodate both the draft model and a safety buffer ($N_{\text{free}}>N_{\text{draft}}+\tau_{\text{low}}$). This hysteresis design ensures that speculative decoding remains disabled during high-load phases (where $N_{\text{free}}<\tau_{\text{low}}$), thereby preventing resource thrashing.
@@ 6.2 Mechanism of KV Cache Expansion
Upon triggering expansion, the system dynamically reallocates GPU memory. The Model Executor allocates a contiguous memory segment of size $S_{\text{draft}}$ to the KV cache pool. Subsequently, the Block Manager updates its internal state: (1) The set of allocatable physical block indices is expanded to include the new range $K_{\text{boundary}},K_{\text{total}})$; (2) Reference counters for these newly instantiated blocks are initialized to zero; (3) The new block indices are appended to the free block queue. Crucially, this process is implemented via in-place expansion, eliminating the need for massive memory copying and minimizing reconfiguration overhead.
@@ 6.3 KV Cache Contraction and Logical Remapping
Figure 7: Overview of the KV Cache Block Contraction and Logical Table Remapping process. The system ensures data consistency during physical memory compaction through vectorized data migration.
@@
Step 1: Identification of Eviction Candidates. The system scans the page tables of all active sequences to identify the set of physical blocks $\mathcal{B}_{\text{evict}}$ located in the memory region to be reclaimed (i.e., blocks with $\text{ID}\geq K_{\text{boundary}}$). These blocks must be migrated to the preserved memory region $0,K_{\text{boundary}})$.
Step 2: Construction of Migration Plan. The system verifies that the free list contains a sufficient number of slots whose physical block IDs are strictly less than $K_{\text{boundary}}$. A one-to-one mapping $\mathcal{M}:b_{\text{old}}\rightarrow b_{\text{new}}$ is then constructed, assigning each block in $\mathcal{B}_{\text{evict}}$ to an available slot in the preserved region $0,K_{\text{boundary}})$. This mapping serves as the basis for the subsequent metadata update and data migration phases.
Step 3: Metadata Update and Remapping. For each block requiring migration, the system performs atomic metadata updates: (a) A new physical block object is instantiated, inheriting the token sequence and size of the original; (b) Predecessor pointers are updated, if a predecessor block is also part of the migration set, the pointer is redirected to its new location; otherwise, the original reference is maintained; (c) The logical Block Table is updated to map the sequence to the new physical block ID; (d) The reference counter for the new block is incremented, and it is marked as allocated. Finally, the old physical block is released, ensuring the logical sequence remains consistent with the physical memory state.
Step 4: Vectorized Data Migration. Following the metadata update, the physical data transfer is executed according to the mapping $\mathcal{M}$. The migration is powered by a custom high-performance kernel (implemented in Triton) utilizing a tile-based parallelization strategy. Each thread block manages the transfer of a complete KV cache block via vectorized load/store instructions. This approach provides robust performance even on hardware lacking advanced virtual memory management (VMM) features like cuMemMap. By decoupling memory reclamation from model execution, we eliminate deadlock risks associated with ”hold-and-wait” patterns.
Step 5: Allocator State Finalization. Post-migration, the block allocator trims its index set, removing IDs $\geq K_{\text{boundary}}$ and clearing associated reference counts, effectively finalizing the contraction.
@@ 6.4 Non-blocking Asynchronous Migration
To minimize latency, we leverage the fact that speculative decoding is inactive during the unloading phase. Consequently, draft model weights are evicted from the GPU immediately, with a backup retained in host memory. Conversely, when memory availability indicates a low request rate, the draft model is reloaded to resume accelerated generation. This reloading process utilizes asynchronous CUDA streams with non-blocking Direct Memory Access (DMA). This allows data transfer to overlap with ongoing computation, ensuring that the background management of model weights incurs negligible impact on the serving latency of the primary model.
@@ 7.2 Dynamic request load
For 7B model under dynamic request rate (Figure 10), as shown in Figure 11, During low request load periods, Nightjar achieves the highest peak throughput by dynamically selecting the optimal speculative length. Conversely, during high-load phases, it strategically disables speculation to revert to standard vanilla decoding. This prevents the severe performance degradation that plagues other speculative methods, allowing Nightjar to maintain consistently superior overall throughput across varying traffic conditions. DSD’s predictions are less accurate during low request load, leading to throughput degradation, but perform better under high request load, while TETRIS exhibits performance drops under high request load due to its high overhead in large batch size and static BanditSpec is not suitable for dynamic batch size.
@@
As summarized in Table IV, Nightjar demonstrates superior performance across all evaluated scenarios. Compared to standard auto-regressive decoding (w/o SD), our method yields a substantial average throughput increase of 27.29%. More importantly, when compared against highly optimized speculative decoding baselines, Nightjar still maintains a clear edge. Specifically, it surpasses DSD and BanditSpec by 22.89% and 19.76% on average, respectively. Even against the strongest baselines, TETRIS and SD, Nightjar achieves competitive average improvements of 7.39% and 8.32%, proving the effectiveness of our proposed design.
Compared to the vanilla auto-regressive baseline (w/o SD), Nightjar achieves an average end-to-end latency reduction of 12.90%. Notably, the performance gains are highly correlated with the target model size. On the 7B model, the latency remains comparable to the baseline, primarily because the overhead of draft token generation and verification offsets the speculative benefits when the base generation speed is already high. However, on the 13B model, where generation costs are significantly heavier, Nightjar demonstrates exceptional efficiency, reducing end-to-end latency by up to 38.35% (on SpecBench).